OpenStudy (anonymous):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation? .

OpenStudy (anonymous):

a b c d are the links someone pls help im stuck on this question

OpenStudy (texaschic101):

need to see the graph

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

@william.1233 can u help me

OpenStudy (texaschic101):

ok....take your 2 points and find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (8,61)...x1 = 8 and y1 = 61 (4,33)...x2 = 4 and y2 = 33 now we sub slope(m) = (33 - 61) / (4 - 8) slope(m) = - 28/-4 slope(m) = 7 y = mx + b slope(m) = 7 use either of your 2 points..(8,61)...x = 8 and y = 61 now sub and solve for b, the y intercept 61 = 7(8) + b 61 = 56 + b 61 - 56 = b 5 = b so your equation for this graph is : y = 7x + 5 (the last link is the answer)

OpenStudy (anonymous):

need to see the graph

OpenStudy (anonymous):

thanks @texaschic101

OpenStudy (texaschic101):

the graph,its been posted...its after the links no problem zlamar..glad to help